# A comprehensive theory for relativistic polaritonic chemistry: a way toward a four component ab initio treatment of molecular systems

### v1

@@ Abstract
We present a new ab initio approach to study molecules containing heavy atoms strongly interacting with quantum fields in optical devices. The relativistic quantum electrodynamics (QED) theory has been rewritten with a formalism close to relativistic quantum chemistry. This general framework represents the ideal starting point to extend the main quantum chemistry methods to relativistic QED. The Polaritonic Dirac Hartree Fock (Pol-DHF) approach is the first method we propose based on this theory. Pol-DHF allows for the simulation of field induced effects on the ground and excited state properties of heavy transition metals molecular complexes. The method is able to include not only the effects of the photons but can be easily extended also to include explicit interactions with positrons. Application of Pol-DHF to three metal hydrides revealed the importance of including radiative QED corrections to the treatment in strong coupling conditions. Due to an accurate description of spin-orbit coupling, the method is able to reproduce polaritonic effects happening at the crossing between singlet and triplet potential energy surfaces.
@@ I Introduction
When matter strongly couples to photons, new hybrid states (polaritons), having partial light and partial matter character, are formed. The strong coupling condition is usually reached inside properly designed optical devices. The simplest example is the Fabry-Pérot cavity [6] (Fig 1), made of two highly reflective planar mirrors, that confine the photons in a small quantization volume and leading to a significant enhancement of the light-matter coupling.
@@
The field properties and therefore the polaritons properties can be controlled changing the geometry and the materials of the cavity. Several demonstrations of polaritonic effects on different physical properties such as absorption spectra, photochemical reaction rates, and conductivity have already been observed in the experiments [2, 7, 3, 5, 8, 9, 10].
Probably the most famous demonstration was obtained, in the experiments performed by the Ebbesen’s group in Strasbourg. In particular, they demonstrated that strong coupling to molecular vibrations can be used to catalyze, slow down or even induce selectivity in chemical reactions [11, 5, 12]. These observations opened a new field that is now known as polaritonic chemistry [13].
@@
In these experiments, the photonic states are usually coupled either to electronic or vibrational states of the molecular systems. However, the electromagnetic nature of the field allows also for modifications of their magnetic properties if a coupling to the spin states is exploited. In this way, a fine control of the magnetizability and aromatic properties of molecules [14], of spin qubits [15, 16] and of spin phases of materials [17, 18] can be obtained. However, reaching the strong coupling condition in this frequency range is impossible using a simple Fabry-Pérot cavity that, in this case, would require a spacing of centimeters between the mirrors. The problem can be circumvented using planar superconducting devices commonly used in Circuit-QED experiments [19] . Using similar devices Affronte’s group has been able to manipulate the spin properties of a Single-Molecule Magnet [16, 20]. Such an accomplishment opened large possibilities of application in spin qubits based quantum computation.
@@
Despite the many improvements in the fabrication of more effective optical devices and the impressive accuracy reached [21, 22] by polaritonic chemistry experiments, many fundamental aspects still remain to be understood. In this context, theory represents a fundamental tool to gain insights on the underlying physics of these processes. In recent years many ab initio methods, able to treat electron-electron and electron-photon correlation at different level of accuracy have been developed [23, 24, 25, 26, 27, 28]. However, they have been mainly applied to investigate cavity induced effects on the electronic and vibrational degrees of freedom of molecular systems. Only very recently, Barlini et al. proposed the first Hartree-Fock based approach to study photon induced effects on the electronic and nuclear magnetic properties of a molecular system [14].
@@
Spin and consequently magnetic properties are intrinsic relativistic features of molecules. An accurate investigation of these properties requires then inclusion of relativistic effects at different level of accuracy. This becomes particularly crucial when, like in Affronte’s experiments [16, 20] Single-Molecule Magnets containing lanthanides atoms need to be used to have sufficiently long living magnetizations.
@@
Moreover, as already discussed in Refs [28, 29] for Van-der Waals interactions, the field has sometimes to power to enhance small effects that are usually neglected in absence of the photons. From these considerations rises the emergency of a consistent relativistic quantum electrodynamical ab initio theory able to include all the necessary effects. Formulating such a general approach is one of the goals of this paper. Similar intent was already presented by Ruggenthaler et al. [23] in a Density Functional Theory (DFT) framework, but only the non-relativistic limit of the method was actually turned into a usable quantum chemistry implementation. Very recently, Konecny et al. [30] proposed a relativistic response theory to investigate electronic excitations of relativistic molecules in optical cavities. Despite this last approach was able to capture interesting effects like cavity-induced singlet-triplet interactions, it is neglecting additional photon driven effects that might be crucial if magnetic properties want to be investigated. Moreover, this approach does not account for modifications induced by the field on the system’s ground state. This is a crucial aspect if we are interested to cavity driven effects on the core properties of relativistic atoms.
@@
In this paper, we develop the first fully relativistic ab initio method to describe the ground state of molecular systems coupled to photons in optical cavities. We start the development from the most general Quantum Electrodynamical Lagrangian formulation to then propose a new and totally method independent approach to develop relativistic ab initio theories for integrating electron-photon systems. This formalism represents the starting point for the development of mean-field but also correlated methodologies. This strategy allows for a more controlled and conscious applications of approximations.
@@
In the current state-of-the-art of relativistic quantum chemistry, the main QED corrections include electron-electron retardation effects (i.e. Breit), and additional radiative QED effects (i.e. vacuum polarization [31, 32], self-energy [33]) responsible for the Lamb-Shift [34]. These additional contributions are divergent and require regularization schemes (e.g. renormalization of the electron mass [35, 36]). Their inclusion in many-electrons systems is usually carried out using effective QED potentials.[37, 38, 39, 40, 41]. In a recent paper [42], it has been demonstrated that Breit contribution and these QED corrections effects on the Ionization Potential (IP) and Electron Affinities (EA) of gold are of the same order of magnitude (0.01-0.03 eV). In our case, instead, interactions with real cavity photons are comparable ( eV for the IP) [43] or (for heavy atoms) smaller than these radiative QED corrections. This clearly prevent us from neglecting a priori these effects. This is a crucial point that will be discussed in this paper and that motivates the proposed general reformulation of the relativistic polaritonic framework.
@@
The paper is structured as follows: in section II a generic derivation of a relativistic QED Hamiltonian theory will be presented, starting from the standard QED Lagrangian. In this section, the choice of the gauge is discussed in detail and the different energy contributions are analyzed. At this stage, no approximations are used in order to set up a theory that could be the starting point for every quantum chemistry modelization. In subsection II.2.2 the coulomb gauge Hamiltonian, after application of the dipole approximation, is used to develop the first Relativistic Polaritonic HF (Pol-DHF) approach. In this context, large space has been dedicated to strategies used to deal with negative energy states and to improve convergence issues. In section III Pol-DHF has been applied to investigate electronic properties of small diatomic molecules containing heavy atoms. We end the paper with conclusions and perspectives.
@@ II Theory
In this section, we follow the formal derivation of relativistic QED theory usually presented in physics text books[44, 45] to develop a Hamiltonian formalism that can be applied to formulate new ab initio methodologies for the simulation of polaritonic molecular systems. This approach allows to keep track of all the approximations used during the derivation, consequently facilitating future improvement strategies, as well as the overall comprehension of the method. The proposed methodology is then used to develop the first Hartree-Fock (HF) based approach for relativistic molecular systems strongly coupled to quantum fields. For convenience reasons, Gaussian units will be used through the whole derivation unless specified otherwise.
@@ II.1 The Quantum Electrodynamics Lagrangian
where and are the vector potential and current associated to the field generated by the nuclei. The field in this last contribution to Eq. 8 due to the Born-Oppenheimer approximation is treated as static quantities. The 4-current in Eq. 8 can be expressed as a function of and as:
@@ II.2 Hamiltonian formulation in Coulomb gauge
From the previous equation one notices the following relationship when :
@@
Substituting Eq.s 25 and 26 in Eq. II.2 we can obtain the following expression for the classical Hamiltonian:
II.2.1 Retardation effects and other QED corrections
@@
In relativistic quantum chemistry, a lot of efforts have been devoted to find effective strategies to include radiative QED corrections [49, 50, 51] (retardation effects, vacuum polarization, Lamb-Shift, etc.), in particular with Bound State QED techniques (BSQED) [52]. Even though such contributions to the electronic properties of the system are fairly small and are in general neglected when performing standard relativistic quantum chemistry calculations, they become quite sizeable when compared with polaritonic effects and should therefore be included.
@@
QED corrections are due to the electromagnetic field-matter interaction term , where . This interaction term contains all the Feynmann graphs related to the various QED corrections.
@@
Such graphs involve what are called virtual-photons, they are internal photonic lines in these graphs and do not correspond to a real particle that can be detected. However, in the formalism of QFT, these virtual photons are the one mediating the electromagnetic interaction between particles. On the other hand, the same interaction term account for the contribution of processes involving real photons. In this case, they appear as external lines in the Feynmann graphs, and they correspond to the particles which are detected in experiments. In BSQED, one is in general interested in the graphs involving virtual photons, whereas in polaritonic chemistry the focus is on the graphs involving real photons. In relativistic polaritonic chemistry, since we cannot a priori neglect radiative QED effects, we need a strategy to include such effects. One strategy could be to adopt a formalism similar to the one of BSQED, however, such formalism, while providing remarkable accuracy for atomics systems, is not really applicable to molecular electronic structure calculations. Fortunately, it is possible to include lower order QED radiative corrections using effective potentials (Breit contribution, Uehling [32, 53, 45, 54], Wichmann-Kroll [55, 54] potential and Self-Energy contribution [37, 38, 45, 54]). Hence, in this way, the dominant graphs involving virtual photons are accounted for, and the formalism of QED is used to describe processes involving real-photons. This strategy remains reasonable since it exploits the linear nature of the Dirac-Hamiltonian and of the interaction term, preserving the additivity of the energy contributions. We would like to remind the reader that some of these contributions include virtual electron-positron pairs. Therefore, using strategies based on effective potentials commonly used in relativistic quantum chemistry are not applicable beyond the no-pair approximation. [54, 56].
@@
Among the radiative QED corrections, the largest term is the so-called Breit term, which is related to retardation effects [53, 45, 57]. Physically, this term originates from the time needed to mediate the electromagnetic interaction between particles, and as the velocity of such particles get closer to the speed of light, such effect becomes more significant. Even though such term contributes to minor energy differences, it can be quite significant for phenomena involving excitation from the core-orbitals, and the effect of such a contribution to the electronic molecular orbital energies has been investigated by Aucar and coworkers [58]. In this work, the possible competition between radiative QED corrections (in particular the Breit contribution) and polaritonic effect is investigated (see Result section III).
@@
There are many ways to derive the Breit term, but one simple way is to work within the Lorenz gauge () which has the advantage of yielding explicitly retarded scalar and vector potentials as done in Ref.s [45]. resulting in an expression of the Full-Breit term as:
@@
which is the retarded electromagnetic interaction without any approximation. For an explicit derivation of this term we refer to Ref.s [53, 45, 54] Since this energy contribution is frequency dependent, it is in general expanded around to obtain the full frequency independent Breit term. Therefore, in relativistic quantum chemistry one generally uses the following 2-body Full-Breit potential:
@@ II.2.2 Hamiltonian Quantization
In the confined space of the optical cavity the wave vector assumes discrete values consequently defining a discrete spectrum of field modes characterized by the direction of the vector and by the polarization of the field oscillations (). In the present case, the discrete vector potential only accounts for real photons, the other contributions due to the transverse part of vector potential (Gaunt, Breit, etc.) could be included as effective potentials (or more elaborate methods based on QED Hamiltonians [52, 49, 50]). In particular, in this work, the Gaunt and Breit contribution are included in the 2-body potential in Eq. 30. Since the molecular system we aim to investigate are usually significantly smaller than the wavelength of the cavity field, we are entitled to apply the dipole approximation imposing that and consequently that in Eq. II.2.2 is equal to 1. This will significantly simplify the photonic part of the Hamiltonian.
Also for the fermionic part, the spinor fields are promoted to spinor fields operators satisfying the following equal-time anti-commutation relations:
@@ II.2.3 Relativistic Pauli-Fierz Hamiltonian in the length gauge
When we deal with molecular systems it is usually more convenient to apply a unitary transformation to the field modes, allowing for a direct coupling between the field operators and the molecular dipole. This transformation take the name of length-gauge transformation:
@@
followed by a rotation of the photonic coordinates . Transformation in Eq.38 will induce a cancellation of the term due to the change in the momentum. The light-matter coupling is now related to the molecular dipole. For details regarding the applications of these transformations, refer to Appendix A and to Ref. [59]. The final Hamiltonian in length gauge appears as:
This Hamiltonian has now a shape that is equivalent to the standard Pauli-Fierz Hamiltonian usually applied in non-relativistic polaritonic chemistry, where molecular orbitals are replaced by molecular spinors. Equation II.2.3 has an apparent origin dependence, coming from the presence of the dipole operator . This problem can be solved by the coherent state transformation with [24]. Applying this transformation to the RPF Hamiltonian, one then can use the coherent state basis for the photonic states and is left with the following Hamiltonian:
In Eq. II.2.3 the presence of the expectation value of the total dipole moments ensures that the Hamiltonian will not explicitly depend on the origin of the reference system. In atomic and molecular physics/chemistry, it is usually more convenient to expand the electronic field on a basis of atomic/molecular orbitals which are solutions of the Dirac equation in an external potential (the Coulomb potential of the nuclei). The field operator expanded on such a basis writes as:
@@
Using these operators the energy components of Hamiltonian II.2.3 can be expressed in a second quantized form implementable in a quantum chemistry code. In the following, we will use a shorthand notation for the mono and bi electronic integrals that is defined as follows:
@@
It is important to notice that Hamiltonian 46 allows both for positive and negative energy states. This point will be discussed in details in the next section (II.2.4).

### v2

# A comprehensive theory for relativistic polaritonic chemistry: a four components ab initio treatment of molecular systems coupled to quantum fields

@@ Abstract
We present a new ab initio approach to study molecules containing heavy atoms strongly interacting with quantum fields in optical devices. The relativistic quantum electrodynamics (QED) theory has been rewritten with a formalism close to relativistic quantum chemistry. This general framework represents the ideal starting point to extend the main quantum chemistry methods to relativistic polaritonic. The Polaritonic Dirac Hartree Fock (Pol-DHF) approach is the first method we propose based on this theory. Pol-DHF allows for the simulation of field induced effects on the ground and excited state properties of heavy transition metals molecular complexes. The method is able to include not only the effects of the photons but can be easily extended also to include explicit interactions with positrons. Application of Pol-DHF to three metal hydrides revealed the importance of including relativistic effects to the treatment in strong coupling conditions. Due to an accurate description of spin-orbit coupling, the method is able to reproduce polaritonic effects happening at the crossing between singlet and triplet potential energy surfaces.
@@ I Introduction
When matter strongly couples to photons, new hybrid states (polaritons), having partial light and partial matter character, are formed. The strong coupling condition is usually reached inside properly designed optical devices. The simplest example is the Fabry-Pérot cavity [6] (Fig 1), made of two highly reflective planar mirrors, that confine the photons, leading to a significant enhancement of the light-matter coupling.
@@
The cavity frequency, and therefore, the polaritons properties can be controlled changing the geometry and the materials in the device. Several manifestations of polaritonic effects on different physical properties such as absorption spectra, photochemical reaction rates, and conductivity have already been observed in the experiments [2, 7, 3, 5, 8, 9, 10]. What is probably the most well-known demonstration of such effects was obtained in the experiments performed by Ebbesen’s group in Strasbourg. In particular, they evidenced that strong coupling to molecular vibrations can be used to catalyze, slow down or even induce selectivity in chemical reactions [11, 5, 12].
These observations opened a new field that is now known as polaritonic chemistry [13].
@@
In these experiments, the photonic states are usually coupled either to electronic or vibrational states of the molecular systems. However, the electromagnetic nature of the field also allows for modifications of their magnetic properties if one exploits the coupling to spin-states. In this way, a fine control of the magnetizability and aromatic properties of molecules [14], of spin qubits [15, 16] and of spin phases of materials [17, 18] can be obtained. However, reaching the strong coupling condition in this frequency range is impossible using a simple Fabry-Pérot cavity that, in this case, would require a spacing of centimeters between the mirrors. The problem can be circumvented using planar superconducting devices commonly used in Circuit-QED experiments [19]. Using similar devices, Affronte’s group has been able to manipulate the spin properties of a Single-Molecule Magnet [16, 20]. Such an accomplishment unlocked many potential applications in spin qubits based quantum computation.
@@
Despite the many improvements in the fabrication of more effective optical devices and the impressive accuracy reached [21, 22] by polaritonic chemistry experiments, many fundamental aspects still remain to be understood. In this context, theory represents a fundamental tool to gain insights on the underlying physics of these processes. In recent years many ab initio methods, able to treat electron-electron and electron-photon correlation at different level of accuracy have been developed [23, 24, 25, 26, 27, 28]. However, they have mainly been applied to investigate cavity induced effects on the electronic and vibrational degrees of freedom of molecular systems. Only very recently, Barlini et al. proposed the first Hartree-Fock based approach to study photon induced effects on the electronic and nuclear magnetic properties of a molecular system [14].
@@
Spin and consequently magnetic properties are intrinsic relativistic features of molecules. An accurate investigation of these properties then requires the inclusion of relativistic effects at different level of accuracy. This becomes particularly crucial when, like in Affronte’s experiments [16, 20, 29, 30, 31, 32, 33] Single-Molecule Magnets containing lanthanides atoms need to be used to have sufficiently long lasting magnetizations.
@@
Moreover, as already discussed in Refs [28, 34] for Van-der Waals interactions, the field can sometimes enhance small effects that are usually negligible in the absence of photons. From these considerations rises the emergency of a consistent relativistic quantum electrodynamical ab initio theory able to include all the necessary effects. Formulating such a general approach is one of the goals of this paper. Similar intent was already presented by Ruggenthaler et al. [23] in a Density Functional Theory (DFT) framework, but only the non-relativistic limit of the method was actually turned into a usable quantum chemistry implementation. Very recently, Konecny et al. [35] proposed a relativistic response theory to investigate electronic excitations of relativistic molecules in optical cavities. Despite this last approach being able to capture interesting effects like cavity-induced singlet-triplet interactions, it is neglecting additional photon driven effects that might be crucial if magnetic properties want to be investigated. Moreover, at the moment, the implemented approach does not account for modifications induced by the field on the system’s ground state. This is a crucial aspect if we are interested to cavity driven effects on the core properties of relativistic atoms.
@@
In this paper, we develop the first fully relativistic wave-function based ab initio method to describe the ground state of molecular systems coupled to photons in optical cavities. We start the development from the most general Quantum Electrodynamics Lagrangian to then propose a new and totally method independent approach to develop relativistic ab initio theories for integrating electron-photon systems. This formalism represents the starting point for the development of mean-field but also correlated methodologies. This strategy allows for a more controlled and conscious applications of approximations.
@@
In the current state-of-the-art of relativistic quantum chemistry, the main QED corrections include electron-electron retardation effects (i.e. Breit), and additional radiative QED effects (i.e. vacuum polarization [36, 37], self-energy [38]) responsible for the Lamb-Shift [39]. These additional contributions are divergent and require regularization schemes (e.g. renormalization of the electron mass [40, 41]). Their inclusion in many-electrons systems is usually carried out using effective QED potentials [42, 43, 44, 45, 46], however, strong efforts are being deployed to accurately include this corrections using the Bound State QED formalism [47, 48, 49]. In a recent paper [50], it has been demonstrated that Breit contribution and these QED corrections effects on the Ionization Potential (IP) and Electron Affinities (EA) of gold are of the same order of magnitude (0.01-0.03 eV). In our case, instead, interactions with real cavity photons are comparable ( eV for the IP) [51] or (for heavy atoms) smaller than these radiative QED corrections. This clearly prevent us from neglecting a priori these effects. This is a crucial point that will be discussed in this paper and that motivates the proposed general reformulation of the relativistic polaritonic framework.
@@
The paper is structured as follows: in section II, a generic derivation of a relativistic QED Hamiltonian theory will be presented, starting from the standard QED Lagrangian. In this section, the choice of the gauge is discussed and the different energy contributions are analyzed. At this stage, no approximations are used in order to set up a theory that could be the starting point for every quantum chemistry modelization. In subsection II.2.2, the coulomb gauge Hamiltonian, after application of the dipole approximation, is used to develop the first Relativistic Polaritonic HF (Pol-DHF) approach. In this context, large space has been dedicated to strategies used to deal with negative energy states and to improve convergence issues. In section III, Pol-DHF has been applied to investigate electronic properties of small diatomic molecules containing heavy atoms. We end the paper with conclusions and perspectives.
@@ II Theory
In this section, we follow the formal derivation of relativistic QED theory usually presented in physics text books[52, 53] to develop a Hamiltonian formalism that can be applied to formulate new ab initio methodologies for the simulation of polaritonic molecular systems. This approach allows to keep track of all the approximations used during the derivation, consequently facilitating future improvement strategies, as well as the overall comprehension of the method. The proposed methodology is then used to develop the first Hartree-Fock (HF) based approach for relativistic molecular systems strongly coupled to quantum fields. For convenience reasons, Gaussian units will be used during the whole derivation unless specified otherwise.
@@ II.1 The Quantum Electrodynamics Lagrangian
where is an external potential associated to the nuclei in the Born-Oppenheimer approximation. The 4-current in Eq. 8 can be expressed as a function of and as:
@@ II.2 Hamiltonian formulation in Coulomb gauge
From the previous equation, one notices the following relationship when :
@@
Substituting Eq.s 25 and 26 in Eq. II.2 we obtain the following expression for the classical Hamiltonian:
II.2.1 Retardation effects and QED corrections
@@
In relativistic quantum chemistry, a lot of efforts have been devoted to find effective strategies to include additional corrections including retardation effects and radiative QED effects (vacuum polarization, Lamb-Shift, etc.) [48, 57, 49], in particular with Bound State QED techniques (BSQED) [47]. Even though such contributions to the electronic properties of the system are fairly small and are in general neglected when performing standard relativistic quantum chemistry calculations, they become quite sizeable when compared with polaritonic effects and should therefore be included.
@@
These additional corrections are due to the electromagnetic field-matter interaction term , where . This interaction term is the one responsible for all these corrections.
@@
In terms of Feynmann diagrams, the latter are related with graphs involving virtual-photons, internal photonic lines in these graphs that do not correspond to a real particle that can be detected. However, in the formalism of QFT, these virtual photons are the one mediating the electromagnetic interaction between particles. On the other hand, the same interaction term account for the contribution of processes involving real photons. In this case, they appear as external lines in the Feynmann graphs, and they correspond to the particles which are detected in experiments. In BSQED, one is in general interested in the graphs involving virtual photons, whereas in polaritonic chemistry the focus is on the graphs involving photons as external lines. In relativistic polaritonic chemistry, since we cannot a priori neglect these additional corrections, we need a strategy to include them into the treatment. One strategy could be adopting a formalism similar to the one used in BSQED. However, such formalism, while providing remarkable accuracy for atomics systems, is not really applicable to molecular electronic structure calculations. In fact, in polaritonic chemistry people mostly rely on the one-mode approximation. In practice, the same approximation will be made in Relativistic Polaritonic chemistry. Most of the effects due to the continuum of modes can be approximated including lower order corrections using effective potentials (Breit contribution, Uehling [37, 58, 53, 59], Wichmann-Kroll [60, 59] potential and Self-Energy contributions [42, 43, 53, 59]). In this way, the dominant graphs involving virtual photons are included while the QED formalism is used to describe processes involving photons as external lines in the diagrams. This strategy remains reasonable since it exploits the linear nature of the Dirac-Hamiltonian and of the interaction term, preserving the additivity of the energy contributions. We would like to remind the reader that some of these contributions include virtual electron-positron pairs and that effective potentials are not applicable beyond the no-pair approximation [59, 61].
@@
Among all aforementioned corrections, we will focus on the so-called Breit term, which is related to retardation effects [58, 53, 62] and in some cases have a magnitude comparable to the other QED corrections effects [50]. Physically, this term originates from the time needed to mediate the electromagnetic interaction between particles, and as the velocity of such particles get closer to the speed of light, such effect becomes more significant. Even though such term contributes to minor energy differences, it can be quite significant for phenomena involving excitation from the core-orbitals, and the effect of such a contribution to the electronic molecular orbital energies has been investigated by Aucar and coworkers [63]. In this work, the possible competition between the aforementioned additional corrections (in particular the Breit contribution) and polaritonic effect is investigated (see Result section III).
@@
There are many ways to derive the Breit term, but one simple way is to work within the Lorenz gauge () which has the advantage of yielding explicitly retarded scalar and vector potentials as done in Ref.s [53] resulting in an expression of the full-Breit term as:
@@
which is the retarded electromagnetic interaction without any approximation. For an explicit derivation of this term we refer to Ref.s [58, 53, 59]. This frequency dependent energy contribution is expanded around to obtain the full frequency independent Breit term. In relativistic quantum chemistry the following 2-body full-Breit potential:
is usually applied.
@@ II.2.2 Hamiltonian Quantization
In the confined space of the optical cavity the wave vector assumes discrete values consequently defining a discrete spectrum of field modes characterized by the direction of the vector and by the polarization of the field oscillations (). In the present case, the discrete vector potential only accounts for contributions due to the transverse part of vector potential (Gaunt, Breit, etc.) could be included as effective potentials (or more elaborate methods based on QED Hamiltonians [47, 48, 57]). In particular, in this work, the Gaunt and Breit contributions are included in the 2-body potential in Eq. 30 at the DHF level of theory using the no-pair approximation. The usual dressed mass of the electron has been used instead of the bare one. Since the molecular systems of interest are usually significantly smaller than the wavelength of the cavity field, we are entitled to apply the dipole approximation imposing that and consequently that in Eq. II.2.2 is equal to 1. This will significantly simplify the photonic part of the Hamiltonian.
Furthermore, for the fermionic part, the spinor fields are promoted to spinor fields operators satisfying the following equal-time anti-commutation relations:
@@ II.2.3 Relativistic Pauli-Fierz Hamiltonian in the length gauge
When investigating molecular systems, it is usually more convenient to apply a unitary transformation to the field modes, allowing for a direct coupling between the field operators and the molecular dipole. This transformation is known as the length-gauge transformation:
@@
followed by a rotation of the photonic coordinates . Transformation in Eq.38 induces a cancellation of the term due to the change in the momentum. The light-matter coupling is now related to the molecular dipole. For details regarding the applications of these transformations, refer to Appendix A and to Ref. [64]. The final Hamiltonian in length gauge appears as:
where and , have been relabeled to emphasize that, due to the length gauge transformation, the latters are modified, and does not coincide with the number of photons anymore [65]. This Hamiltonian is the analogous of the standard Pauli-Fierz Hamiltonian usually applied in non-relativistic polaritonic chemistry [23, 66, 24], where molecular orbitals are replaced by molecular spinors. Equation II.2.3 has an apparent origin dependence, coming from the presence of the dipole operator . This problem can be solved by the coherent state transformation with [24]. Applying this transformation to the Relativistic-Pauli-Fiertz (RPF) Hamiltonian, one then can use the coherent state basis for the photonic states and is left with the following Hamiltonian:
where is the usual one-electron Dirac Hamiltonian. In Eq. II.2.3, the presence of the expectation value of the total dipole moments () ensures that the Hamiltonian will not explicitly depend on the origin of the reference system. In atomic and molecular physics/chemistry, it is usually more convenient to expand the electronic field on a basis of atomic/molecular orbitals which are solutions of the Dirac equation in an external potential (the Coulomb potential of the nuclei). The field operator expanded on such a basis writes as:
@@
Using these operators the energy components of Hamiltonian II.2.3 can be expressed in a second quantized form implementable in a quantum chemistry code. In the following, we will use a shorthand notation for the mono- and bi- electronic integrals that are defined as follows:
@@
where refers to the molecular quadrupole and to the polariton number. Notice that Hamiltonian 46 allows both for positive and negative energy states. This point will be discussed in details in the next section (II.2.4).